OpenStudy (anonymous):

Solve for x and y using elimination. −2x−5y=−19 x+2y=9 *Please help me out with the plugging in of the numbers into x and y! :)*

OpenStudy (anonymous):

multiply x+2y=9 by 2 then add them together

OpenStudy (anonymous):

So I get 2x+4y=18?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now add the first equation with the new one

OpenStudy (anonymous):

-3y=-10?

OpenStudy (anonymous):

−2x−5y=−19 2x+4y=18 -y=-1 y=1

OpenStudy (anonymous):

AH! My bad! I had all the math right I just solved it weird! OK so if y=1 then how do I solve for x?

OpenStudy (anonymous):

now substitute the value of y in one of the equations to get x

OpenStudy (anonymous):

x=7?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

omg your my lifesaver no joke thank you so much!!!!!!!!!!!

OpenStudy (anonymous):

any time